There are 6 boys and 6 girls in the finals of a talent contest. A contest is held to pick the top 3 winners in both the boy and girl groups in order of talent. How many different options for winners are there?
Answer by greenestamps(13200)



For both boys and girls, the number of ways to rank the top 3 out of 6 is

P(6,3) = 6*5*4 = 120

So the number of ways of ranking the top 3 of both boys and girls is

120*120 = 14400

ANSWER: 14400